OpenStudy (anonymous):

The force of gravity varies inversely with the square of the distance between given planet and the sun because \[F_g=\frac{GMm}{r^2}\]

OpenStudy (anonymous):

@Jemurray3

OpenStudy (anonymous):

what is the "varies" symbol again? \[F\;\;\; \frac 1{r^2}\]

OpenStudy (anonymous):

\[ \propto \] \propto which is short for "proportional to"

OpenStudy (anonymous):

"varies inversely" means proportional to \(\frac 1 r\) correct? \(\frac 1 {r^2}\)

OpenStudy (anonymous):

if A varies inversely with B, you say \[ A \propto \frac{1}{B} \] here, you're saying "The force (A) is varies inversely with the square of the distance (B)". So A is the force, and B is r^2. \[ F \propto \frac{1}{r^2} \]